# Low-Bandwidth Matrix Multiplication: Faster Algorithms and More General Forms of Sparsity

### v1

@@ Abstract
In prior work, Gupta et al. (SPAA 2022) presented a distributed algorithm for multiplying sparse $n\times n$ matrices, using $n$ computers. They assumed that the input matrices are uniformly sparse—there are at most $d$ non-zeros in each row and column—and the task is to compute a uniformly sparse part of the product matrix. Initially each computer knows one row of each input matrix, and eventually each computer needs to know one row of the product matrix. In each communication round each computer can send and receive one $O(\log n)$-bit message. Their algorithm solves this task in $O(d^{1.907})$ rounds, while the trivial bound is $O(d^{2})$.
@@
We show that we can still compute $X=AB$ in $O(d^{1.832})$ rounds even if one of the three matrices ($A$, $B$, or $X$) is average-sparse instead of uniformly sparse. We present algorithms that handle a much broader range of sparsity in $O(d^{2}+\log n)$ rounds, and present conditional hardness results that put limits on further improvements and generalizations.
@@ 1 Introduction
In this work we study sparse matrix multiplication in a distributed setting, for the low-bandwidth model (which is closely related to the node-capacitated clique model). Our work improves over the prior work by [12] in two ways:
  1. 1.
    we present a faster algorithm, and
  2. 2.
    we study more general forms of sparsity.
@@ 1.2 Contribution 1: faster algorithm
The algorithm from prior work [12] is based on the idea of processing triangles. Here a triangle is a triple $\{i,j,k\}$ such that $A_{ij}$ and $B_{jk}$ are nonzeros and $X_{ik}$ is one of the elements of interest, and we say that we process the triangle if we add the product $A_{ij}B_{jk}$ to the sum $X_{ik}$. Once all such triangles are processed, we have computed all elements of interest in the product matrix.
There are two strategies that [12] uses for processing triangles:
@@ 1.3 Contribution 2: beyond uniform sparsity
Perhaps the most natural candidate to consider would be average sparsity: we would simply assume that the total number of nonzeros in $A$ and $B$ and the total number of elements of interest in $X$ is at most $dn$. Unfortunately, it turns out that this notion of sparsity is way too general, and we cannot expect fast algorithms for this case: in Lemma 6.3 we show that if we could solve average-sparse matrix multiplication for $n\times n$ matrices using $n$ computers in time that is independent of $n$ or only mildly depends on $n$ (say, polylogarithmic in $n$), it would imply major breakthroughs in our understanding of dense matrix multiplication.
@@
Now equipped with this notation, we can rephrase the main result from the prior work [12]: they show that $[\mathsf{US}:\mathsf{US}:\mathsf{US}]$ can be solved in $O(d^{1.927})$ rounds for semirings and in $O(d^{1.907})$ rounds for fields. On the other hand, in Lemma 6.3 we argue why it is unlikely to extend this all the way to $[\mathsf{AS}:\mathsf{AS}:\mathsf{AS}]$. The key question is what happens between the two extremes, $\mathsf{US}$ and $\mathsf{AS}$? In particular, could we solve $[\mathsf{BD}:\mathsf{BD}:\mathsf{BD}]$ efficiently?
Our answer to this question consists of four parts:
@@
    We generalize prior work and show that we can tolerate one average-sparse matrix: $[\mathsf{US}:\mathsf{US}:\mathsf{AS}]$ can be computed in $O(d^{1.867})$ rounds for semirings and in $O(d^{1.832})$ rounds for fields. We emphasize that here the running time is still independent of $n$.
@@
    As soon as two of the matrices are $\mathsf{BD}$ or more general, the task becomes at least as hard as computing the sum of values distributed across $n$ computers, or broadcasting a single value to $n$ computers. Hence, it is natural to expect an additional $O(\log n)$ term in the running time.
@@
    We show that many combinations all the way to $[\mathsf{BD}:\mathsf{AS}:\mathsf{AS}]$ and $[\mathsf{US}:\mathsf{AS}:\mathsf{GM}]$ can be solved in $O(d^{2}+\log n)$ rounds, for both semirings and fields.
@@
    Finally, we provide evidence that this is essentially as general as what we can hope for.
The results are summarized in Table 2.
Table 2: Summary of results (for semirings)
Sparsity Upper bound Conditional
hardness
$[\mathsf{US}:\mathsf{US}:\mathsf{US}]$ $O(d^{1.927})$ [12]
$[\mathsf{US}:\mathsf{US}:\mathsf{US}]$ $O(d^{1.867})$ Theorem 4.2
@@
$[\mathsf{US}:\mathsf{US}:\mathsf{GM}]$ $O(d^{2}+\log n)$ Theorem 5.3
$[\mathsf{US}:\mathsf{BD}:\mathsf{BD}]$ $O(d^{2}+\log n)$ Theorem 5.3 Lemma 6.1
@@
$[\mathsf{BD}:\mathsf{BD}:\mathsf{BD}]$ $O(d^{2}+\log n)$ Theorem 5.11 Lemma 6.1
@@
$[\mathsf{US}:\mathsf{GM}:\mathsf{GM}]$ Lemma 6.5
@@
$[\mathsf{BD}:\mathsf{BD}:\mathsf{GM}]$ Lemma 6.7
@@
$[\mathsf{AS}:\mathsf{AS}:\mathsf{AS}]$ Lemma 6.3
@@ 1.4 Related work and applications
Similar to [12], we work in the supported version of the low-bandwidth model. In general, the supported model [11, 10, 18] refers to a setting in which the structure of the input is known in advance (we can do arbitrary preprocessing based on the structure of the input), while the specific instance is revealed at run time. In the case of graph problems, we can see an unweighted graph $G$ in advance, while the adversary reveals a weighted version $G^{\prime}$ of $G$ (a special case being a $0/1$-weighted graph, which can be interpreted as a subgraph of $G$). For matrix multiplication, we know the sparsity structure in advance (i.e., which elements are potentially nonzero), while the adversary reveals the concrete values at run time.
@@ 1.5 Open questions for future work
Our hardness results are conditional. Turning these into unconditional lower bounds is yet another question for future work.
Finally, there are some gaps in Table 2. For example, we do not have conditional hardness result for $[\mathsf{US}:\mathsf{US}:\mathsf{GM}]$, and also Lemmas 6.5 and 6.7 do not cover all permutations of the matrix families.
@@ 2 Preliminaries
For our algorithms, it does not matter how the input and output is distributed among the computers—with an additional $O(d)$ time we can permute the input and output as appropriate. Our conditional lower bounds hold for any fixed distribution of input and output.
@@ 3 Handling few triangles fast
We will start by a general result that will form the foundation for all of our upper-bound results. [12, Lemma 5.2] observed that if we have a balanced instance in which each node touches at most $\kappa$ triangles, we can solve matrix multiplication in $O(\kappa)$ rounds. However, the key challenge is what to do in unbalanced instances: we have at most $\kappa n$ triangles, but some nodes can touch many more than $\kappa$ triangles. The following lemma shows how to handle also such instances efficiently:
@@
We emphasize that this strictly improves over [12, Lemma 5.1] in two different ways: our result is applicable in a more general setting, and we avoid the factor-$2$ loss in the exponent of the running time.
@@ 4.2 Part 1: handling dense parts
This is a generalization of [12, Lemma 3.1], which holds for $[\mathsf{US}:\mathsf{US}:\mathsf{US}]$. The original proof makes use of the following facts about $G(\mathcal{T})$:
@@
Then by plugging in Lemma 4.7 instead of [12, Lemma 3.1] in the proof of [12, Lemma 4.1], we obtain the following corollary:
@@
We follow the strategy in the proof of [12, Lemma 4.1]: Apply Lemma 4.7 to find a cluster $U$. Put the triangles $T\in\mathcal{T}[U]$ that are fully contained in $U$ to $\mathcal{P}$, and put the triangles $T\in\mathcal{T}$ that only partially touch $U$ to $\mathcal{T}^{\prime}$. Repeat until there are sufficiently few triangles left; put all remaining triangles to $\mathcal{T}^{\prime}$.
@@
Finally, by plugging in Lemma 4.9 instead of [12, Lemma 4.1] in the proof of [12, Lemma 4.2], we obtain the following corollary:
@@
We follow the proof of [12, Lemma 4.2]: Apply Lemma 4.9 repeatedly to construct clusterings $\mathcal{P}_{1},\mathcal{P}_{2},\dotsc$. Stop once the residual set of triangles $\mathcal{T}^{\prime}$ is sufficiently small. The original analysis holds verbatim.
@@ 5.2 Algorithm for $[\mathsf{BD}:\mathsf{AS}:\mathsf{AS}]$
6 Conditional hardness
In this section we connect the hardness of sparse matrix multiplication with simpler primitives, such as broadcasting.
@@ 6.1 Broadcasting and aggregation
For settings between $[\mathsf{US}:\mathsf{BD}:\mathsf{BD}]$ and $[\mathsf{US}:\mathsf{AS}:\mathsf{GM}]$ our upper bound from Theorem 5.3 has an additive $O(\log n)$ term, which comes from broadcast and convergecast operations. In the following lemma, we show that some broadcasting and/or aggregation is needed in all of these cases. We note that a simple low-bandwidth adaptation of the proof in [8] shows an $\Omega(\log n)$-round lower bound for computing a sum over values held by all computers, and hence this result also implies an unconditional lower bound.
@@
Now this is equivalent to computing a sum $x=\sum_{j}a_{j}$, in a setting in which initially each computer holds one $a_{j}$.

### v2

@@ Abstract
In prior work, Gupta et al. (SPAA 2022) presented a distributed algorithm for multiplying sparse $n\times n$ matrices, using $n$ computers. They assumed that the input matrices are uniformly sparse—there are at most $d$ non-zeros in each row and column—and the task is to compute a uniformly sparse part of the product matrix. The sparsity structure is globally known in advance (this is the supported setting). As input, each computer receives one row of each input matrix, and each computer needs to output one row of the product matrix. In each communication round each computer can send and receive one $O(\log n)$-bit message. Their algorithm solves this task in $O(d^{1.907})$ rounds, while the trivial bound is $O(d^{2})$.
@@
We present a near-complete classification of the complexity of matrix multiplication for all combinations of these notions of sparsity. We show that almost all cases fall in one of these classes:
  1. 1.
    We have an upper bound of $O(d^{1.832})$ rounds for computing $X=AB$. An example is the case where $X$ and $A$ are uniformly sparse but $B$ is average-sparse; this is a generalization of the prior work beyond the uniformly-sparse case.
  2. 2.
    We have a lower bound of $\Omega(\log n)$ and an upper bound of $O(d^{2}+\log n)$ rounds for computing $X=AB$. An example is the case where $X$, $A$, and $B$ have bounded degeneracy.
  3. 3.
    We have a lower bound of $n^{\Omega(1)}$ rounds for computing $X=AB$. An example is the case where $X$ and $A$ have bounded degeneracy but $B$ is a general matrix.
  4. 4.
    We have a conditional lower bound: a fast algorithm for computing $X=AB$ would imply major improvements in dense matrix multiplication. An example is the case where $X$, $A$, and $B$ are average-sparse matrices.
Our work highlights the role that bounded degeneracy has in the context of distributed matrix multiplication: it is a natural class of sparse matrices that is restrictive enough to admit very fast algorithms—much faster than what we can expect for the average-sparse case.
@@ 1 Introduction
In this work we study sparse matrix multiplication in a distributed setting, for the low-bandwidth model (which is closely related to the node-capacitated clique model). Our work improves over the prior work by [13] in two ways: we present a faster algorithm, and we study more general forms of sparsity.
@@ 1.2 Contribution 1: faster algorithm
The algorithm from prior work [13] is based on the idea of processing triangles. Here a triangle is a triple $\{i,j,k\}$ such that $A_{ij}$ and $B_{jk}$ are nonzeros and $X_{ik}$ is one of the elements of interest, and we say that we process the triangle if we add the product $A_{ij}B_{jk}$ to the sum $X_{ik}$. Once all such triangles are processed, we have computed all elements of interest in the product matrix. There are two strategies that [13] uses for processing triangles:
@@ 1.3 Contribution 2: beyond uniform sparsity
Perhaps the most natural candidate to consider would be average sparsity: we would simply assume that the total number of nonzeros in $A$ and $B$ and the total number of elements of interest in $X$ is at most $dn$. Unfortunately, it turns out that this notion of sparsity is way too general, and we cannot expect fast algorithms for this case: in Theorem 6.19 we show that if we could solve average-sparse matrix multiplication for $n\times n$ matrices using $n$ computers in time that is independent of $n$ or only mildly depends on $n$ (say, polylogarithmic in $n$), it would imply major breakthroughs in our understanding of dense matrix multiplication.
@@
Now equipped with this notation, we can rephrase the main result from the prior work [13]: they show that $[\mathsf{US}:\mathsf{US}:\mathsf{US}]$ can be solved in $O(d^{1.927})$ rounds for semirings and in $O(d^{1.907})$ rounds for fields. On the other hand, in Theorem 6.19 we argue why it is unlikely to extend this all the way to $[\mathsf{AS}:\mathsf{AS}:\mathsf{AS}]$. The key question is what happens between the two extremes, $\mathsf{US}$ and $\mathsf{AS}$? In particular, could we solve $[\mathsf{BD}:\mathsf{BD}:\mathsf{BD}]$ efficiently?
In this work we present a near-complete classification of the complexity of matrix multiplication for all possible combinations of classes $\mathsf{US}$, $\mathsf{BD}$, $\mathsf{AS}$, and $\mathsf{GM}$. We show that almost all cases fall in one of these classes:
@@
    We have an upper bound of $O(d^{1.832})$ rounds by Theorem 4.2. For example, $[\mathsf{US}:\mathsf{US}:\mathsf{AS}]$ falls in this class.
@@
    We have a lower bound of $\Omega(\log n)$ rounds by Theorem 6.15 and an upper bound of $O(d^{2}+\log n)$ rounds by Theorems 5.3 and 5.11. For example, $[\mathsf{BD}:\mathsf{BD}:\mathsf{BD}]$ falls in this class.
@@
    We have a lower bound of $\Omega(\sqrt{n})$ rounds by Theorem 6.27. For example, $[\mathsf{BD}:\mathsf{BD}:\mathsf{GM}]$ falls in this class.
@@
    We have a conditional lower bound by Theorem 6.19: a fast algorithm would imply major improvements in dense matrix multiplication. For example, $[\mathsf{AS}:\mathsf{AS}:\mathsf{AS}]$ falls in this class.
There is one outlier in our current classification: $[\mathsf{US}:\mathsf{US}:\mathsf{GM}]$ admits a trivial algorithm that runs in $O(d^{4})$ rounds, but we do not know if it can be solved in $O(d^{1.832})$ rounds.
Our results are summarized in Table 2. The table summarizes the results for the case of semirings, but analogous results hold for the case of fields—the only difference is that the exponent $1.867$ is replaced with $1.832$ and we have got $1.1571\leq\lambda\leq 2-2/\omega<$ 1.157.
Table 2: Summary of results (for semirings). Here $31\leq\lambda\leq 4/$ 3 is the infimum of exponents $\lambda$ such that dense matrix multiplication for semirings can be done in $O(n^{\lambda})$ rounds. The lower bounds marked with $\dagger$ only hold for certain permutations of the matrix families.
Sparsity Upper bound Lower bound
$[\mathsf{US}:\mathsf{US}:\mathsf{US}]$ $O(d^{1.867})$ Theorem 4.2 $\Omega(d^{\lambda})$ trivial
@@
$[\mathsf{US}:\mathsf{US}:\mathsf{GM}]$ $O(d^{4})$ trivial $\Omega(d^{\lambda})$ trivial
$[\mathsf{US}:\mathsf{BD}:\mathsf{BD}]$ $O(d^{2}+\log n)$ Theorem 5.3 $\Omega(d^{\lambda})$, $\Omega(\log n)$ Theorem 6.15
@@
$[\mathsf{BD}:\mathsf{BD}:\mathsf{BD}]$ $O(d^{2}+\log n)$ Theorem 5.11 $\Omega(d^{\lambda})$, $\Omega(\log n)$ Theorem 6.15
@@
$[\mathsf{US}:\mathsf{GM}:\mathsf{GM}]$ $\Omega(\sqrt{n})$ Theorem 6.27
@@
$[\mathsf{BD}:\mathsf{BD}:\mathsf{GM}]$ $\Omega(\sqrt{n})$ Theorem 6.27
@@
$[\mathsf{AS}:\mathsf{AS}:\mathsf{AS}]$ $\Omega(n^{(\lambda-1)/2})$ Theorem 6.19
@@
1.4 Key conceptual message: role of bounded degeneracy
Our work highlights the role that the class of bounded-degeneracy matrices plays in the theory and practice of distributed matrix multiplication. While the class of average-sparse matrices is arguably more natural, it is also too broad to admit fast algorithms, even if we are in the supported setting (Theorem 6.19). On the other hand, the class of uniformly-sparse matrices familiar from prior work can be far too restrictive for many applications. Our work suggests a promising new notion of sparsity in this context: matrices with bounded degeneracy. First, such matrices naturally arise especially in graph-theoretic application, and they also contain as special cases the important classes of row-sparse and column-sparse matrices. Second, our work shows that bounded-degeneracy matrices admit very efficient algorithms in the low-bandwidth setting (and as a corollary, also in any setting that is at least as strong as this model), at least if we know the structure of the matrices in advance (Theorem 5.11). We believe this can open a new research direction that studies the complexity of matrix multiplication across different distributed and parallel settings for bounded-degeneracy matrices.
@@ 1.5 Related work and applications
Similar to [13], we work in the supported version of the low-bandwidth model. In general, the supported model [12, 11, 21] refers to a setting in which the structure of the input is known in advance (we can do arbitrary preprocessing based on the structure of the input), while the specific instance is revealed at run time. In the case of graph problems, we can see an unweighted graph $G$ in advance, while the adversary reveals a weighted version $G^{\prime}$ of $G$ (a special case being a $0/1$-weighted or $1/\infty$-weighted graph, which can be interpreted as a subgraph of $G$). For matrix multiplication, we know the sparsity structure in advance (i.e., which elements are potentially nonzero), while the adversary reveals the concrete values at run time.
Our logarithmic lower bounds build on the classic results in the CREW PRAM model by Cook, Dwork and Reischuk [8] and Dietzfelbinger, Kutylowski and Reischuk [9]. As observed by e.g. Roughgarden, Vassilvitskii and Wang [20], such lower bounds can be applied also in more modern settings.
@@ 1.6 Open questions for future work
Finally, there are some gaps in Table 2. For example, we do not know if $[\mathsf{US}:\mathsf{US}:\mathsf{GM}]$ is strictly harder than $[\mathsf{US}:\mathsf{US}:\mathsf{US}]$, and also Theorem 6.27 do not cover all permutations of the matrix families. Finally, we do not know if $O(d^{2}+\log n)$ can be further improved to, say, $O(d^{1.867}+\log n)$.
@@ 2 Preliminaries
For our algorithms, it does not matter how the input and output is distributed among the computers—with an additional $O(d)$ time we can permute the input and output as appropriate. Our lower bounds hold for any fixed distribution of input and output that may also depend on the support.
@@ 3 Handling few triangles fast
We will start by a general result that will form the foundation for all of our upper-bound results. [13, Lemma 5.2] observed that if we have a balanced instance in which each node touches at most $\kappa$ triangles, we can solve matrix multiplication in $O(\kappa)$ rounds. However, the key challenge is what to do in unbalanced instances: we have at most $\kappa n$ triangles, but some nodes can touch many more than $\kappa$ triangles. The following lemma shows how to handle also such instances efficiently:
@@
We emphasize that this strictly improves over [13, Lemma 5.1] in two different ways: our result is applicable in a more general setting, and we avoid the factor-$2$ loss in the exponent of the running time.
@@ 4.2 Part 1: handling dense parts
This is a generalization of [13, Lemma 3.1], which holds for $[\mathsf{US}:\mathsf{US}:\mathsf{US}]$. The original proof makes use of the following facts about $G(\mathcal{T})$:
@@
Then by plugging in Lemma 4.7 instead of [13, Lemma 3.1] in the proof of [13, Lemma 4.1], we obtain the following corollary:
@@
We follow the strategy in the proof of [13, Lemma 4.1]: Apply Lemma 4.7 to find a cluster $U$. Put the triangles $T\in\mathcal{T}[U]$ that are fully contained in $U$ to $\mathcal{P}$, and put the triangles $T\in\mathcal{T}$ that only partially touch $U$ to $\mathcal{T}^{\prime}$. Repeat until there are sufficiently few triangles left; put all remaining triangles to $\mathcal{T}^{\prime}$.
@@
Finally, by plugging in Lemma 4.9 instead of [13, Lemma 4.1] in the proof of [13, Lemma 4.2], we obtain the following corollary:
@@
We follow the proof of [13, Lemma 4.2]: Apply Lemma 4.9 repeatedly to construct clusterings $\mathcal{P}_{1},\mathcal{P}_{2},\dotsc$. Stop once the residual set of triangles $\mathcal{T}^{\prime}$ is sufficiently small. The original analysis holds verbatim.
@@ 5.2 Algorithm for $[\mathsf{BD}:\mathsf{AS}:\mathsf{AS}]$
6 Lower bounds
Our lower bounds are organized as follows:
  • Section 6.1 presents lower bounds that build on the hardness of broadcasting and aggregation. For example, an extreme case of $\mathsf{BD}\times\mathsf{BD}=\mathsf{US}$ is the multiplication of one dense row with one dense column. This enables us to compute arbitrary dot products, and in particular arbitrary sums. Hence matrix multiplication is at least as hard as computing the sum $n$ values. Similar arguments show that matrix multiplication is at least as hard as broadcasting one value to $n$ computers. We show that both of these tasks require $\Omega(\log n)$ rounds.
  • Section 6.2 presents lower bounds that connect sparse matrix multiplication with dense matrix multiplication. If we have an algorithm for computing $\mathsf{AS}\times\mathsf{AS}=\mathsf{AS}$, we can pack a small dense matrix into one corner, and this way solve an arbitrary dense matrix multiplication. Particular care is needed here to account for the dual role that $n$ plays in our setting: it is both the matrix dimension and the number of computers.
  • Section 6.3 presents lower bounds that arise from communication complexity arguments. The key observation is that sparse matrix multiplication can be used to transmit information between two parties, Alice holding a subset of computers (and hence is able to manipulate their input) while Bob holds the rest of the computers (and hence is able to see their output). If there is a fast algorithm for matrix multiplication, Alice and Bob can simulate it to exchange information with each others too efficiently, leading to a contradiction.
Here it is good to note that the lower bounds in Section 6.1 are fundamentally related to the number of messages (they hold even if the computers can send in each round one message that is arbitrarily large), while the lower bounds in Section 6.3 are fundamentally related to the total size of messages.
@@ 6.1 Broadcasting and aggregation
For settings between $[\mathsf{US}:\mathsf{BD}:\mathsf{BD}]$ and $[\mathsf{US}:\mathsf{AS}:\mathsf{GM}]$ our upper bound from Theorem 5.3 has an additive $O(\log n)$ term, which comes from broadcast and convergecast operations. In the following lemma, we show that some broadcasting and/or aggregation is indeed needed in all of these cases.
@@
Now this is equivalent to computing a sum $x=\sum_{j}a_{j}$, in a setting in which initially each computer holds one $a_{j}$ and one computer has to hold $x$.